CO was allowed to react with Fe2O3 in the amount of 43.95grams.what is the amount of Fe that will be created because of the reaction?


1
Expert's answer
2021-10-19T23:04:01-0400

The mass of iron oxide which reacts with 36.03 g of carbon = 159.69 g


The mass of iron oxide which reacts with 8.320 g of carbon= (159.69 x8.320) / 32.03 = 41.48 g


2.


2Na + Cl_2 \to 2NaCl


According to the chemical equation written above


2 moles of Na (23 x 2 = 46 g ) reacts with amount of chlorine = 1 mol (70.9 g)


7.589 g of Na reacts with amount of chlorine = (70.9 x 7.589) / 46 = 11.697 g


Given mass of chlorine = 30.54 g


Na is limiting reagent as it will decide the yield of the reaction products.


and Cl2 is excess reagent.


Now , theoretical yield of NaCl can be calculated as :


46 g of sodium reacts with chlorine to give = 58.44 g of NaCl


7.589 g of Na reacts with chlorine and will give = (58.44 x 7.589) / 46 =9.641 g


Theoretical yield = 9.641 g


Excess reagent used = 11.697 g


Excess reagent leftover= 30.540 g - 11.697 g =`8.843 g


Actual yield = 1.635 g


Yield % =( Actual yield x 100) / theoretical yield = (1.635 x 100) / 9.641 = 16.95 %